# Non-standard analysis for coherent risk estimation: hyperfinite representations, discrete Kusuoka formulae, and plug-in asymptotics

### v1

@@ 9. Hyperfinite bootstrap validity
We establish bootstrap validity by exploiting the internal nature of the resampling procedure. In the hyperfinite framework, the bootstrap distribution is simply the internal distribution of the statistic conditional on the hyperfinite sample.
@@ 9.1. Internal bootstrap construction
9.2. Near-standardness of conditional laws
Bootstrap validity requires that the conditional law of $\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})$ is infinitely close to the limiting law of $\sqrt{N}(\hat{\rho}_{N}-\rho(X))$.
Assumption 9.2 (Bootstrap regularity).
The random variable $X$ satisfies:
  1. (B1)
    $F_{X}$ is continuous;
  2. (B2)
    $X$ has a density $f_{X}$ that is bounded and bounded away from 0 on the interior quantile range $\{q_{X}(\alpha):\alpha\in[\delta,1-\delta]\}$ for each fixed $\delta\in(0,1/2)$.
Equivalently, (B2) says $f_{X}$ is continuous and strictly positive on every compact subinterval of the interior of the support of $X$. Under (B1)–(B2), the quantile map $F\mapsto q_{F}$ is Hadamard differentiable at $F_{X}$, and the spectral functional $T(F)=-\int_{0}^{1}q_{F}(\alpha)\varphi(\alpha)\,d\alpha$ inherits Hadamard differentiability; see [18, Lemma 21.3, Thm. 21.5].
Lemma 9.3 (Hadamard derivative of the spectral functional).
Proof.
$T(F)=-\int_{0}^{1}Q(F)(\alpha)\varphi(\alpha)\,d\alpha,$
the chain rule gives
Theorem 9.4 (Hyperfinite Bootstrap Consistency).
Let $X\in L^{2}$ satisfy Assumption 9.2, and let $\varphi$ be Lipschitz. Assume moreover that the hypotheses of Theorem 10.5 hold, so that the asymptotic variance $\sigma_{\varphi}^{2}$ in (28) is finite. Let $\mathcal{L}^{*}(\cdot|X)$ denote the internal conditional law given the sample (i.e., the law under the bootstrap-index randomness $(U_{i})$ with $(X_{i})$ fixed). Then, for $L({}^{*}\mathsf{P})$-almost all sample paths $\omega$,
@@ 9.2. Near-standardness of conditional laws
$d_{K}\left(\mathcal{L}^{*}\left(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\;\Big|\;X(\omega)\right),\ N(0,\sigma_{\varphi}^{2})\right)\approx 0,$
where $\sigma_{\varphi}^{2}$ is the asymptotic variance in Theorem 10.5.
Proof.
The proof combines the standard functional delta method for the bootstrap as explained by [18, Thm. 23.9] with NSA reasoning, using the internal Kolmogorov distance $d_{K,m}$ from Lemma 9.1 to ensure measurability.
Near-standardness of the empirical measure. Fix a sample path $\omega$ such that the empirical CDF $\hat{F}_{N}$ satisfies $\|\hat{F}_{N}-F_{X}\|_{\infty}\approx 0$. By the hyperfinite Glivenko–Cantelli theorem (Theorem 3.20), this holds for $L({}^{*}\mathsf{P})$-almost all $\omega$.
Standard bootstrap delta method. The standard bootstrap delta method [18, Thm. 23.9] asserts: under Assumption 9.2, the conditional distribution of $\sqrt{n}(T(\hat{F}_{n}^{*})-T(\hat{F}_{n}))$ given the sample converges to $\mathcal{L}(T^{\prime}_{F_{X}}(\mathbb{G}))$ in probability in the bounded-Lipschitz metric, where $\mathbb{G}$ is an $F_{X}$-Brownian bridge. Since the limit distribution $N(0,\sigma_{\varphi}^{2})$ has a continuous CDF, this implies convergence in Kolmogorov distance [18, Lemma 21.2]: for every $\varepsilon,\delta>0$ and $m\in\mathbb{N}$, there exists $n_{0}\in\mathbb{N}$ such that for all $n\geqslant n_{0}$,
$\mathsf{P}\Big(d_{K,m}\big(\mathcal{L}(\sqrt{n}(T(\hat{F}_{n}^{*})-T(\hat{F}_{n}))\mid X),\,N(0,\sigma_{\varphi}^{2})\big)>\varepsilon\Big)<\delta.$
NSA repackaging via internal distance. Fix standard $\varepsilon,\delta>0$ and standard $m\in\mathbb{N}$. The event $E_{m,\varepsilon}:=\{d_{K,m}(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),N(0,\sigma_{\varphi}^{2}))>\varepsilon\}$ is internal by Lemma 9.1(ii), hence Loeb measurable.
The sample size $N$ exceeds any standard threshold $n_{0}$ (since $N$ is infinite), so $^{*}\mathsf{P}(E_{m,\varepsilon})<\delta$. Taking Loeb measures and using that $\delta>0$ is arbitrary, $L({}^{*}\mathsf{P})(E_{m,\varepsilon})=0$ for each standard $\varepsilon>0$ and $m\in\mathbb{N}$.
Intersecting over countably many $\varepsilon=1/k$ and $m\in\mathbb{N}$, we obtain a Loeb-probability-one set $G$ on which $d_{K,m}(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),N(0,\sigma_{\varphi}^{2}))<1/r$ for all standard $m,r\in\mathbb{N}$. To obtain an infinite $M$ with $d_{K,M}\approx 0$, we use saturation: for each standard $r\in\mathbb{N}$, define the internal set
$A_{r}:=\{m\in{}^{*}\mathbb{N}:d_{K,m}(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),N(0,\sigma_{\varphi}^{2}))<1/r\}.$
@@
On the set $G$, each $A_{r}$ contains all standard $m$. The family $\{A_{r}\}_{r\in\mathbb{N}}\cup\{B_{s}\}_{s\in\mathbb{N}}$ (where $B_{s}:=\{m\in{}^{*}\mathbb{N}:m>s\}$) has the finite intersection property. By countable saturation, there exists $M\in\bigcap_{r\in\mathbb{N}}A_{r}\cap\bigcap_{s\in\mathbb{N}}B_{s}$, which is infinite (since $M>s$ for all standard $s$) and satisfies $d_{K,M}<1/r$ for all standard $r$, hence $d_{K,M}\approx 0$. By Lemma 9.1(i), the standard-part distribution equals $N(0,\sigma_{\varphi}^{2})$.
Identification of the asymptotic variance. The random variable $T^{\prime}_{F_{X}}(\mathbb{G})$ equals
$\int_{0}^{1}\frac{\mathbb{G}(q_{X}(\alpha))}{f_{X}(q_{X}(\alpha))}\,\varphi(\alpha)\,d\alpha=\int_{0}^{1}\frac{B(\alpha)}{f_{X}(q_{X}(\alpha))}\,\varphi(\alpha)\,d\alpha,$
where $B(\alpha):=\mathbb{G}(q_{X}(\alpha))$ is a standard Brownian bridge on $[0,1]$ (since $\mathbb{G}(x)$ at $x=q_{X}(\alpha)$ with $F_{X}(q_{X}(\alpha))=\alpha$ gives $\mathrm{Cov}(B(\alpha),B(\beta))=\min(\alpha,\beta)-\alpha\beta$).
This integral is Gaussian with mean zero. Under the stronger smoothness hypotheses of Theorem 10.5 (in particular, $q_{X}\in C^{1}$ with $q_{X}^{\prime}=1/f_{X}(q_{X})$), the variance equals
$\sigma_{\varphi}^{2}=\int_{0}^{1}\int_{0}^{1}(\min(\alpha,\beta)-\alpha\beta)\,\frac{\varphi(\alpha)\varphi(\beta)}{f_{X}(q_{X}(\alpha))f_{X}(q_{X}(\beta))}\,d\alpha\,d\beta.$
Since the limit $N(0,\sigma_{\varphi}^{2})$ has a continuous CDF, weak convergence to it implies Kolmogorov distance convergence (see, e.g., [18, Lemma 21.2]). Thus $d_{K}(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),N(0,\sigma_{\varphi}^{2}))\approx 0$. ∎
Remark 9.5 (The power of the hyperfinite viewpoint).
The advantage of this approach is that we do not need to check “sequences” of conditional measures converging in a double-limit sense. We fix a single infinite $N$ and a typical path; the bootstrap distribution is then a fixed internal distribution, which we show is near-standard to the Gaussian. The transfer principle converts the standard Hadamard differentiability into an internal statement, and near-standardness of the empirical measure (from Theorem 3.20) provides the bridge between the internal and standard worlds.
@@ 10.1. Hyperfinite central limit theorem
where $\Phi$ is the standard normal distribution function.
@@ 10.2. Asymptotic normality of spectral estimators
Under the hypotheses of Theorem 10.5, the asymptotic linearity of the spectral L-statistic can be expressed as follows: for every $\varepsilon\in\mathbb{Q}_{>0}$, there exists $n_{0}\in\mathbb{N}$ such that for all $n\geqslant n_{0}$,
(26) $\mathsf{P}\Big(\Big|\sqrt{n}(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X))-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathrm{IF}_{\varphi}(X_{i})\Big|>\varepsilon\Big)<\varepsilon,$
where $\mathrm{IF}_{\varphi}(x):=\int_{0}^{1}\varphi(\alpha)\,q_{X}^{\prime}(\alpha)\,(\mathbf{1}_{x\leqslant q_{X}(\alpha)}-\alpha)\,d\alpha$. This statement is first-order in the sense that it can be formalised using only bounded quantification over rationals, hence is amenable to transfer and underspill.
@@
This is the content of [17, Thm. 8.5]: under the smoothness hypotheses (bounded $\varphi$, $q_{X}\in C^{1}$ with finite $\int\varphi^{2}(q_{X}^{\prime})^{2}$), the L-statistic $\hat{\rho}_{n,\varphi}$ is asymptotically linear with influence function $\mathrm{IF}_{\varphi}$. The $o_{\mathsf{P}}(1)$” remainder term means precisely that for every $\varepsilon>0$, $\mathsf{P}(|\text{remainder}|>\varepsilon)\to 0$, which can be restated in the $\forall\varepsilon\exists n_{0}\forall n\geqslant n_{0}$” form above. ∎
@@
Let $X\in L^{2}$ and $\varphi$ a bounded spectrum. Assume the quantile function $q_{X}$ is continuously differentiable on $(0,1)$ with derivative $q_{X}^{\prime}=1/f_{X}(q_{X})$, where $f_{X}$ is the density of $X$. Assume moreover that $\int_{0}^{1}\varphi(\alpha)^{2}q_{X}^{\prime}(\alpha)^{2}\,d\alpha<\infty$, so that $\sigma_{\varphi}^{2}$ in (28) is finite. Then
@@
By Lemma 10.4, the asymptotic linearity statement (26) holds for all $n\geqslant n_{0}$ for some standard $n_{0}$. For the fixed infinite $N$ in our standing model, $N>n_{0}$, so
$^{*}\mathsf{P}\Big(\Big|\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))-T_{N}\Big|>\varepsilon\Big)<\varepsilon.$
Intersecting over a sequence $\varepsilon=1/m$ ($m\in\mathbb{N}$) and taking Loeb probabilities, we see that $\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))\approx T_{N}$ holds $L({}^{*}\mathsf{P})$-almost surely.
@@
Transfer back to standard sequences. To obtain the standard convergence in distribution, we use overspill on a complement set. Fix $t\in\mathbb{Q}$ and rational $\varepsilon>0$. Define the internal set
@@
By transfer of (26) with $\varepsilon=1/m$, every infinite $n$ satisfies
@@
Let $B_{t,\varepsilon}:={}^{*}\mathbb{N}\setminus S_{t,\varepsilon}$ be the internal complement. Since $N\in S_{t,\varepsilon}$ for every infinite $N$, the set $B_{t,\varepsilon}$ contains no infinite hypernaturals. If $B_{t,\varepsilon}$ contained arbitrarily large standard integers, then by overspill (Proposition 3.21) it would contain an infinite hypernatural, a contradiction. Hence there exists $n_{0}\in\mathbb{N}$ such that $B_{t,\varepsilon}\subseteq\{1,\dots,n_{0}-1\}$, and therefore every standard $n\geqslant n_{0}$ lies in $S_{t,\varepsilon}$.

### v2

@@ 9. Hyperfinite bootstrap validity
We establish bootstrap validity for canonical spectral plug-in estimators. We first state and prove the result in standard terms, then discuss its NSA reformulation.
@@ 9.1. Internal bootstrap construction
9.2. Bootstrap regularity
Assumption 9.1 (Bootstrap regularity).
The random variable $X$ satisfies:
  1. (B1)
    $F_{X}$ is continuous;
  2. (B2)
    $X$ has a density $f_{X}$ that is continuous and strictly positive on the interior quantile range $\{q_{X}(\alpha):\alpha\in[\delta,1-\delta]\}$ for each fixed $\delta\in(0,1/2)$.
In particular, (B2) implies that $f_{X}$ is bounded and bounded away from 0 on every compact subinterval of the interior of the support of $X$. Under (B1)–(B2), the quantile map $F\mapsto q_{F}$ is Hadamard differentiable at $F_{X}$, and the spectral functional $T(F)=-\int_{0}^{1}q_{F}(\alpha)\varphi(\alpha)\,d\alpha$ inherits Hadamard differentiability; see [18, Lemma 21.3, Thm. 21.5].
Lemma 9.2 (Hadamard derivative of the spectral functional).
Proof.
$T(F)=-\int_{0}^{1}Q(F)(\alpha)\varphi(\alpha)\,d\alpha,$
the chain rule gives
9.3. Bootstrap validity
Theorem 9.3 (Bootstrap validity for the canonical spectral estimator).
Let $X\in L^{2}$ satisfy Assumption 9.1, and let $\varphi$ be Lipschitz. Let $(X_{1}^{*},\dots,X_{n}^{*})$ be an Efron bootstrap sample drawn from $(X_{1},\dots,X_{n})$. Assume moreover that the hypotheses of Theorem 10.5 hold, so that the asymptotic variance $\sigma_{\varphi}^{2}$ in (28) is finite. Then
$d_{K}\!\left(\mathcal{L}\!\left(\sqrt{n}\bigl(\hat{\rho}_{n,\varphi}(X_{1}^{*},\dots,X_{n}^{*})-\hat{\rho}_{n,\varphi}(X_{1},\dots,X_{n})\bigr)\,\Big|\,X_{1},\dots,X_{n}\right),N(0,\sigma_{\varphi}^{2})\right)\xrightarrow{\mathsf{P}}0.$
Proof.
Let
$T(F):=-\int_{0}^{1}q_{F}(\alpha)\varphi(\alpha)\,d\alpha,\qquad T_{n}:=T(\hat{F}_{n}),\qquad T_{n}^{*}:=T(\hat{F}_{n}^{*}).$
Here
$\hat{F}_{n}^{*}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}^{*}},\qquad\mathsf{P}^{*}(\cdot):=\mathsf{P}(\cdot\mid X_{1},\dots,X_{n}).$
Since $\varphi$ is Lipschitz, Lemma 7.5 gives
$|\hat{\rho}_{n,\varphi}(X_{1},\dots,X_{n})-T_{n}|\leqslant\frac{L}{n}\cdot\frac{1}{n}\sum_{i=1}^{n}|X_{i}|,$
and similarly
$|\hat{\rho}_{n,\varphi}(X_{1}^{*},\dots,X_{n}^{*})-T_{n}^{*}|\leqslant\frac{L}{n}\cdot\frac{1}{n}\sum_{i=1}^{n}|X_{i}^{*}|.$
Therefore
$\sqrt{n}\,|\hat{\rho}_{n,\varphi}(X_{1},\dots,X_{n})-T_{n}|\leqslant\frac{L}{\sqrt{n}}\cdot\frac{1}{n}\sum_{i=1}^{n}|X_{i}|\xrightarrow{\mathsf{P}}0$
by the strong law. Conditionally on the sample, Markov’s inequality yields, for every $\eta>0$,
$\mathsf{P}^{*}\!\left(\sqrt{n}\,|\hat{\rho}_{n,\varphi}(X_{1}^{*},\dots,X_{n}^{*})-T_{n}^{*}|>\eta\right)\leqslant\frac{L}{\eta\sqrt{n}}\cdot\frac{1}{n}\sum_{i=1}^{n}|X_{i}|,$
and the right-hand side converges to 0 in $\mathsf{P}$-probability.
By Lemma 9.2, $T$ is Hadamard differentiable at $F_{X}$. Hence the standard bootstrap delta method [18, Thm. 23.9] yields conditional weak convergence in probability (equivalently, convergence in the bounded-Lipschitz metric). Since the limit $N(0,\sigma_{\varphi}^{2})$ has a continuous CDF, [18, Lemma 21.2] upgrades this to convergence in Kolmogorov distance:
$d_{K}\!\left(\mathcal{L}\!\left(\sqrt{n}(T_{n}^{*}-T_{n})\,\big|\,X_{1},\dots,X_{n}\right),N(0,\sigma_{\varphi}^{2})\right)\xrightarrow{\mathsf{P}}0.$
The previously established $\sqrt{n}$-equivalence allows $T_{n}$ and $T_{n}^{*}$ to be replaced by $\hat{\rho}_{n,\varphi}(X_{1},\dots,X_{n})$ and $\hat{\rho}_{n,\varphi}(X_{1}^{*},\dots,X_{n}^{*})$, respectively. ∎
9.4. Internal Kolmogorov distance and NSA reformulation
@@
Remark 9.5 (NSA reformulation of bootstrap validity).
Since $d_{K,m}\leqslant d_{K}$ for every standard $m\in\mathbb{N}$, Theorem 9.3 implies that for each fixed standard $m$,
$d_{K,m}\bigl(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),\,N(0,\sigma_{\varphi}^{2})\bigr)\approx 0$
on a $L({}^{*}\mathsf{P})$-probability-one event. Fix a sample path in this event. For each $r\in\mathbb{N}$, define
@@
$A_{r}:=\left\{m\in{}^{*}\mathbb{N}:\ d_{K,m}\bigl(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),\,N(0,\sigma_{\varphi}^{2})\bigr)<\frac{1}{r}\right\},$
and, for each $s\in\mathbb{N}$,
$B_{s}:=\{m\in{}^{*}\mathbb{N}:\ m>s\}.$
@@
Each $A_{r}$ is internal and contains all standard integers. Therefore the family $\{A_{r}\}_{r\in\mathbb{N}}\cup\{B_{s}\}_{s\in\mathbb{N}}$ has the finite intersection property. By countable saturation, there exists an infinite
$M\in\bigcap_{r\in\mathbb{N}}A_{r}\cap\bigcap_{s\in\mathbb{N}}B_{s}.$
Hence
$d_{K,M}\bigl(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),\,N(0,\sigma_{\varphi}^{2})\bigr)\approx 0,$
and Lemma 9.4(i) shows that the conditional standard-part distribution equals $N(0,\sigma_{\varphi}^{2})$.
@@ 10.1. Hyperfinite central limit theorem
where $\Phi$ is the standard normal distribution function. (When $\sigma=0$ the conclusion is trivial: $S_{N}=0$ for all $N$, and the degenerate distribution $\delta_{0}$ is recovered.)
@@ 10.2. Asymptotic normality of spectral estimators
Under the hypotheses of Theorem 10.5, for every $\varepsilon,\delta\in\mathbb{Q}_{>0}$ there exists $n_{0}\in\mathbb{N}$ such that for all $n\geqslant n_{0}$,
(26) $\mathsf{P}\Bigg(\Bigg|\sqrt{n}\bigl(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X)\bigr)-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathrm{IF}_{\varphi}(X_{i})\Bigg|>\varepsilon\Bigg)<\delta,$
where
$\mathrm{IF}_{\varphi}(x):=\int_{0}^{1}\varphi(\alpha)\,q_{X}^{\prime}(\alpha)\,\bigl(\mathbf{1}_{\{x\leqslant q_{X}(\alpha)\}}-\alpha\bigr)\,d\alpha.$
In particular, the remainder is $o_{\mathsf{P}}(1)$ in a form suitable for transfer.
@@
This is the content of [17, Thm. 8.5]: under the smoothness hypotheses (bounded $\varphi$, $q_{X}\in C^{1}$ with finite $\int\varphi^{2}(q_{X}^{\prime})^{2}$), the L-statistic $\hat{\rho}_{n,\varphi}$ is asymptotically linear with influence function $\mathrm{IF}_{\varphi}$. The $o_{\mathsf{P}}(1)$” remainder term means precisely that for every $\varepsilon>0$ and $\delta>0$, $\mathsf{P}(|\text{remainder}|>\varepsilon)<\delta$ for all sufficiently large $n$, which can be restated in the $\forall\varepsilon,\delta\in\mathbb{Q}_{>0}\;\exists n_{0}\;\forall n\geqslant n_{0}$” form above. ∎
@@
Let $X\in L^{2}$ and $\varphi$ a bounded spectrum. Assume the quantile function $q_{X}$ is continuously differentiable on $(0,1)$ with derivative $q_{X}^{\prime}=1/f_{X}(q_{X})$, where $f_{X}$ is the density of $X$. Assume moreover that $\int_{0}^{1}\varphi(\alpha)^{2}q_{X}^{\prime}(\alpha)^{2}\,d\alpha<\infty$, so that $\sigma_{\varphi}^{2}$ in (28) is finite. If $\sigma_{\varphi}^{2}=0$, the conclusion is that $\sqrt{n}(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X))\xrightarrow{\mathsf{P}}0$; in the sequel we assume $\sigma_{\varphi}^{2}>0$. Then
@@
By Lemma 10.4, for every standard $\eta,\delta>0$,
$^{*}\mathsf{P}\Big(\Big|\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))-T_{N}\Big|>\eta\Big)<\delta.$
Since $\delta>0$ is arbitrary, the Loeb measure of the event in brackets is 0. Thus, for every standard $\eta>0$,
$L({}^{*}\mathsf{P})\Big(\Big|\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))-T_{N}\Big|>\eta\Big)=0.$
Intersecting over $\eta=1/m$, $m\in\mathbb{N}$, we obtain
$\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))\approx T_{N}\qquad\text{$L({}^{*}\mathsf{P})$-almost surely.}$
@@
Transfer back to standard sequences. To obtain the standard convergence in distribution, we use countable saturation on the complement set. Fix $t\in\mathbb{Q}$ and rational $\varepsilon>0$. Define the internal set
@@
By transfer of (26) with $\varepsilon$ replaced by $1/m$ and $\delta$ replaced by $1/m$, every infinite $n$ satisfies
@@
Let $B_{t,\varepsilon}:={}^{*}\mathbb{N}\setminus S_{t,\varepsilon}$. This is an internal subset of $^{*}\mathbb{N}$, and it contains no infinite hypernaturals.
If $B_{t,\varepsilon}\cap\mathbb{N}$ were unbounded, then for each standard $m\in\mathbb{N}$ the internal set
$B_{t,\varepsilon}\cap\{n\in{}^{*}\mathbb{N}:\ n>m\}$
would be non-empty. These sets clearly have the finite intersection property. By countable saturation, their intersection would contain some $H\in{}^{*}\mathbb{N}$ with $H>m$ for every standard $m$, i.e. an infinite hypernatural $H\in B_{t,\varepsilon}$, a contradiction.
Therefore $B_{t,\varepsilon}\cap\mathbb{N}$ is bounded. Hence there exists $n_{0}\in\mathbb{N}$ such that every standard $n\geqslant n_{0}$ lies in $S_{t,\varepsilon}$.